OpenStudy (anonymous):

3x^2-3x^3/x^3-x simplify and state restrictions (rational variable expressions)

OpenStudy (anonymous):



OpenStudy (anonymous):

factor and cancel

OpenStudy (anonymous):

wow that was wrong

OpenStudy (anonymous):

\[\frac{3x^2-3x^3}{x^3-x}\] \[=\frac{3x^2(1-x)}{x(x^2-1)}\] \[=\frac{3x^2(1-x)}{x(x+1)(x-1)}\] there, that's better

OpenStudy (anonymous):

then cancel an \(x\) and also \(\frac{1-x}{x-1}=-1\) making the whole thing \[\frac{-3x}{x+1}\]